OpenStudy (anonymous):

The product of two consecutive positive odd integers is 255. What are the integers?

OpenStudy (anonymous):

\[x \times(x +2)=255\] \[x ^{2}+2x \] \[x ^{2}+2x - 255 = 0\] This has two possible solutions: x = 15 x = 17 Since we know the numbers are odd, the answer is 15. To check this is true: 15 x 17 = 255

OpenStudy (anonymous):

okay thank you :) that is what i got as well